Quit Smoking: Previous studies suggest that use of nicotine-replacement therapies and antidepressants can help people stop smoking. The New England Journal of Medicine published the results of a double-blind, placebo-controlled experiment to study the effect of nicotine patches and the antidepressant bupropion on quitting smoking. The target for quitting smoking was the 8th day of the experiment.

In this experiment researchers randomly assigned smokers to treatments. Of the 162 smokers taking a placebo, 28 stopped smoking by the 8th day. Of the 272 smokers taking only the antidepressant buproprion, 82 stopped smoking by the 8th day.

Calculate the 99% confidence interval to estimate the treatment effect of buproprion (placebo-treatment). (The standard error is about 0.0407. Use critical value z = 2.576.)

Please show all the steps. (in legible handwriting or excel)

ANSWER: (______, ______)
Answer #1

Answer

Given that

p1 = x1/n1 = 28/162

p2 = x2/n2 = 82/272

standard error for proportion difference is 0.0407

z critical = 2.576

Confidence interval=